Exercise 3-s mcwork en March 31 of the oarrest yeor indludes the following selected accounts before adhjusting entries bave been prepared Prepaid Insurance Supplies 27,460 $7.905 19.990 Notes Payable Its by Study Unearned Rent Revenue Rent Reverue Interest Expense Sislaries and Wages Expense 62 420 An analysis of the accounts shows the following 1. The equipment depreciates $263 per month 2. One-third of the unearned rent was earmed as revenue during the quarter 3. Interest of $530 is accrued on the notes payable 4. Supplies on hand total $678 s. Insurance expires at the rate of $271 per month
Prepare the adjusting entries at March 31, assuming that adjusting entes are made urtety Adsioonal account titles are automatically indented when amseount s entered. Do not indent manually. If no entry is requiredt accounts are Depredation Expense, Insurance Expense, Interest Payable, and Supplies Expenses. (Credit aousentered. Do not indent manually. If no entry is requiredd, select No entry for the account titfes and enter 0 for the amousts.) No. Account Titles and Explanation Debit Credit 2. 3. 4 5.
Answer #1

Please find below answer and please provide rating.. let me know if you need any clarification..

No Account title and explanation Dr Cr
1 Depreciation 789
Accumulated depreciation-equipment 789
To record 3 months of depreciation, 263*3 = 789
2 unearned revenue 1150
Rent revenue 1150
To record earning one-third of the rent revenue, 3450/3 = 1150
3
Interest expense 530
Interest payable 530
To record the accrual of interest on the notes payable
4
Supplies expense 1926
Supplies 1926
To record the use of supplies, 2604-678 = 1926
5
Insurance expense 813
Prepaid insurance 813
To record the expiration of insurance, 271*3 = 813